• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 22:52:14 on April 7,2007

    Entry number 198122

    keyword=Target and Magnets stable, keeping CHL 20.5 g/s to target for now.

    Cryo has somewhat remedied the backflow troubles we were experiencing.  Magnets 
    and target appear stable for now.  Below is a capture of their ELOG.  As they 
    state... if problems come back, we may need to dial down our flow (and possibly 
    lower our beam current to compensate for the loss in target cooling).
    


    FIGURE 1